# Hybrid Monte Carlo for Failure Probability Estimation with Gaussian Process Surrogates

### v1

@@ 2.1 Gaussian process surrogates
$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right).$
Returning to the Herbie example, the entropy surface for the trained GP is shown in the right panel of Figure 1 and along the slice in Figure 2 (green dotted/dashed). Notice the high entropy regions mostly agree with the locations of the true failure contour (we could improve upon this with a CL design instead of an LHS). A reasonable acquisition could fall in any of these local optima. Booth et al., (2024b) extended these ideas, with some necessary adaptations, to enable contour location with Bayesian DGP surrogates. Henceforth, all GPs and DGPs we consider will be from sequential CL designs.
@@ 2.2 Importance sampling
$\hat{\alpha}_{\mathrm{IS}}=\frac{1}{B}\sum_{i=1}^{B}w_{i}\;\mathbb{I}_{\left\{f(\mathbf{x}_{i})>t\right\}}\quad\textrm{where}\quad w_{i}=\frac{p(\mathbf{x}_{i})}{q(\mathbf{x}_{i})}\quad\textrm{for}\quad\mathbf{x}_{i}\sim q(\mathbf{x}).$ (6)
The “bias weights” $w_{i}$ account for the difference between the input density and the bias density. This estimator is theoretically unbiased and requires a smaller budget of simulator evaluations ($B\ll M$) if the bias distribution appropriately covers and targets the failure region (Srinivasan,, 2002). But estimation of an effective bias distribution is tricky. The ideal bias distribution is $q(\mathbf{x})^{*}=\frac{1}{\alpha}\mathbb{I}_{\left\{f(\mathbf{x})>t\right\}}p(\mathbf{x})$, but naturally neither $\alpha$ nor the indicator function are known a priori. For the sake of feasibility, $q(\mathbf{x})$ is traditionally constrained to a known distributional family, such as a Gaussian mixture model (GMM; Reynolds et al.,, 2009). Bias distribution training is then relegated to learning distributional parameters (in a GMM, these include the number of mixtures, mixture weights, and component means and variances), often accomplished through cross validation or information criteria maximization.
In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large sample of the input distribution, then train a GMM bias distribution to these predicted failures. If the surrogate effectively identifies the failure region, then these samples should mimic the ideal $q(\mathbf{x})^{*}$. With a trained bias distribution in-hand, samples may be drawn from $q(\mathbf{x})$ and deployed in Eq. (6) to provide an estimate of the failure probability.
While SIIS has been widely embraced for its unbiasedness, it has several key weaknesses. First, the bias distribution, which features heavily in the final estimate of Eq. (6), is a completely auxilary quantity. Hypothetically, two different bias distributions could produce the same samples $\left\{\mathbf{x}_{i}\right\}_{i=1}^{B}$, thus producing the same $\left\{f(\mathbf{x}_{i})\right\}_{i=1}^{B}$. But in the calculation of Eq. (6), these two bias distributions would contribute different weights, resulting in different final answers even though the training data from the computer simulator was exactly the same. O’Hagan, (1987) first pointed out this conundrum as a violation of the “Likelihood principle.” A second weakness is the restriction of $q(\mathbf{x})$ to distributional families like GMMs. Failure regions may be complex and may not fit nicely under these target distributions. Consequently, SIIS typically requires large $B$ in order to adequately cover the failure region. For context, Peherstorfer et al., (2016) entertained $B\in\left\{100,1000,10000,100000\right\}$ and found performance continued to improve as $B$ increased. Furthermore, there is no clear avenue for determining the budget allocation between surrogate training and importance sampling estimation. Finally, the SIIS method does not leverage the full potential of the surrogate model. Many expensive simulator evaluations go into training the surrogate, and in the end the surrogate is only used to identify predicted failures, with no mention of its UQ. We suspect this is a byproduct of Peherstorfer et al., using reduced order models, which do not provide the UQ of a GP, in their seminal work. Ultimately, we propose a flexible estimation procedure which avoids bias distributions altogether, leverages surrogate UQ, and enables strategic budget allocation.
3 Hybrid Monte Carlo with GPs
In this section, we introduce a failure probability estimation scheme which (i) relies only on the surrogate model and the simulator data, (ii) leverages surrogate UQ in the estimation phase, and (iii) can provide effective estimates with a budget of only several hundred. Specifically, we will upgrade the Monte Carlo estimate of Eq. (2) so it requires fewer evaluations of $f(\mathbf{x})$ with minimal reduction in accuracy.
Our approach (and those of our competitors) may be organized into two sequential phases: the training of the surrogate and the estimation of the failure probability. Let $N$ denote the number of simulator evaluations used to train the surrogate, and let $B$ denote the remaining budget of simulator evaluations available for estimation. For the duration of this section, we assume both $N$ and $B$ are pre-determined; in Section 4 we will consider how to allocate evaluations between them. Recall, we are targeting expensive simulations where the total budget, $N+B$, is constrained to be in the hundreds.
3.1 Methodology
Let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive). We may partition the integral of Eq. (1) into two components as follows, with the corresponding Monte Carlo estimate,
@@ 3.1 Methodology
Now, replace expensive evaluation of $f(\mathbf{x}_{i})$ outside of $\mathcal{U}$ with the surrogate prediction $\mu_{N}(\mathbf{x}_{i})$, resulting in the following “hybrid Monte Carlo” estimator:
@@
The hybrid nature of this estimator is in the use of both true simulator evaluations (left term) and surrogate evaluations (right term). Each $\mathbf{x}_{i}$ will fall under either the left indicator term or the right one, never both. The computational cost of this estimator is greatly reduced since evaluations of the surrogate are effectively free. The accuracy of this estimator is contingent upon the classification accuracy of the surrogate and the choice of $\mathcal{U}$. Here, we focus on the choice of $\mathcal{U}$, presuming we already have a trained GP-based surrogate in-hand. [Training a surrogate for contour location has been previously investigated, as we reviewed in Section 2.1.]
@@
Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize our contribution with a discrete set of samples, which form the basis of any Monte Carlo estimator. Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$. To deploy our estimator $\hat{\alpha}_{\textrm{HYB}}$, we need a decision rule for classifying each $\mathbf{x}_{i}$ into or out of $\mathcal{U}$. In a perfect world, $\mathcal{U}$ would capture all the samples which the surrogate would misclassify. We are not privy to this information, but we can leverage the surrogate’s UQ to identify the samples at which it is most uncertain. Recall, entropy (Eq. 5) quantifies the degree of uncertainty in the pass/fail prediction of a surrogate. Accordingly, our decision rule is strikingly simple – place the $B$-many samples from $X_{M}$ with the highest entropy into $\mathcal{U}$. Denote these as $X_{M}^{\mathcal{U}}$. Hopefully our choice of the letter $\mathcal{U}$ is now apparent - it denotes the inputs with the highest “$\mathcal{U}$ncertainty”. Revisiting Figure 2, notice the high entropy regions (green dotted/dashed) align well with the misclassification regions (yellow). The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations.
Li and Xiu, (2010) employed a hybrid estimator similar to that of Eq. (7), but they chose to allocate expensive simulator evaluations to samples predicted to be closest to the contour. We suspect this choice was motivated by their use of polynomial chaos expansion models as surrogates which do not facilitate UQ. Nevertheless, we will implement this allocation scheme with our hybrid Monte Carlo estimator as a competitor in our benchmark exercises later.
For a fixed $N$, $B$, and $M$, our method proceeds as follows. First, train the surrogate using $N$-many evaluations of the high fidelity simulator. Our method is agnostic to surrogate and design choice as long as the surrogate provides posterior predictive distributions for the calculation of entropy, but we prefer contour locating designs with GP or DGP surrogates. Second, sample a very large number of points $X_{M}\sim p(\mathbf{x})$ and use the surrogate to calculate the classification entropy for each sample following Eq. (5). Third, identify the $B$-many points in $X_{M}$ with the highest entropy and evaluate the simulator at these locations. Fourth, update the surrogate with these new observations of the simulator. Finally, estimate $\hat{\alpha}_{\textrm{HYB}}$ following Eq. (7) by combining the simulator evaluations at $X_{M}^{\mathcal{U}}$ with surrogate predictions at all other samples of $X_{M}$.
This fourth step – updating the surrogate given the simulator data at the high-entropy sampled locations – is significant. In the end, the surrogate is trained on $N+B$ observations, making full use of all the available data. Since the $X_{M}^{\mathcal{U}}$ points are likely near a failure contour, they could significanly improve the accuracy of the surrogate near the contour. The ability to update the surrogate and make full use of all simulator data is a unique advantage over SIIS, whose final estimate is contingent only upon the $B$-many points sampled from the bias distribution.

### v2

@@ 2.2 Contour location
$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right).$
Returning to the Herbie example, the entropy surface for the trained GP is shown in the right panel of Figure 1 and along the slice in Figure 2 (green dotted/dashed). Notice the high entropy regions mostly agree with the locations of the true failure contour (we could improve upon this with a CL design instead of an LHS). A reasonable acquisition could fall in any of these local optima.
Naturally, implementations of entropy as an aqcuisition function vary on the finer details. For GP surrogates we will leverage the optimization scheme of Cole et al., (2023), which first deploys a small LHS, identifies the LHS sample with the highest entropy, then runs a numerical optimization of entropy initialized at that point. By avoiding multi-starts, Cole et al., target local optima in the entropy surface. For DGP surrogates, we will follow Booth et al., (2024b). To avoid cumbersome numerical optimizations with the heftier MCMC-based DGP, Booth et al., use triangulation candidates (Gramacy et al.,, 2022), then select the candidate on the Pareto front of entropy and uncertainty. Henceforth, all GPs and DGPs we consider will be trained with these two contour location methods, respectively.
Figure 3: Various designs for estimating the failure probability of the Herbie function with a GP surrogate and a total budget of $N=150$. All methods start with the same CL design of size $n=100$ (black circles), leaving a remaining budget of $B=50$ (blue triangles) for either continued surrogate training or importance sampling. Left: Entire budget is spent on continued contour location, with acquisitions exploring the entire space. Center: Remaining budget is sampled from a bias distribution for importance sampling (black circles omitted since they do not feature in the SIIS estimate). Right: Remaining budget is used for MC samples with highest entropy, solely exploiting near the predicted contour. Comparison of failure probability estimates obtained from these is provided later in Section 4. Although there is no clear distinction in performace between the left/right designs for this simple illustrative example, the continued exploration of the “All CL” design will hinder performance on larger, harder problems (Figure 6).
Continuing the Herbie function example, the left panel of Figure 3 shows the design points chosen from entropy-based contour location with a GP surrogate (Cole et al.,, 2023), indicated by both the black circles and blue triangles (the distinction between the two will become clear when we discuss the other panels). Notice, the chosen points attempt to balance exploration (points spread out over the entire domain) and exploitation (points hovering around the red failure contour). The resulting surrogate may inform an $\hat{\alpha}_{\textrm{SURR}}$ estimate (4); we will include this as a competitor in our later exercises.
@@ 2.3 Importance sampling
The “bias weights” $w_{i}$ account for the difference between the input density and the bias density. This estimator is theoretically unbiased and requires a smaller budget of simulator evaluations ($B\ll M$) if the bias distribution appropriately covers and targets the failure region (Srinivasan,, 2002). But estimation of an effective bias distribution is tricky. The ideal bias distribution is $q(\mathbf{x})^{*}=\frac{1}{\alpha}\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}p(\mathbf{x})$, but neither $\alpha$ nor the indicator function are known a priori. For the sake of feasibility, $q(\mathbf{x})$ is traditionally constrained to a known distributional family, such as a Gaussian mixture model (GMM; Reynolds et al.,, 2009). Bias distribution training is then relegated to learning distributional parameters (in a GMM, these include the number of mixtures, mixture weights, and component means and variances), often accomplished through cross validation or information criteria maximization.
In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large Monte Carlo sample of the input distribution, then train a GMM bias distribution to these predicted failures. If the surrogate effectively identifies the failure region, then these samples should mimic the ideal $q(\mathbf{x})^{*}$. With a trained bias distribution in-hand, samples may be drawn from $q(\mathbf{x})$, evaluated through the expensive simulator, and deployed in Eq. (6) to provide an estimate of the failure probability.
To illustrate, consider our running Herbie function example. Using the GP surrogate trained on the 100-point contour locating design from Figure 3 (black circles) and a large MC sample from the input distribution, we trained a GMM bias distribution to the MC samples which the surrogate predicted to be failures. The blue triangles in the center panel show 50 points sampled from this bias distribution. These samples may be evaluated through the true model to estimate $\hat{\alpha}_{\textrm{IS}}$ (6), thereby exhausting our entire budget of 150 evaluations. While the blue triangles target the failure region, they do not provide enough coverage to ensure a precise estimate. In our later experiments (Section 4), SIIS performance is highly variable.
While SIIS has been widely embraced for its unbiasedness, it has several key weaknesses. First, the bias distribution, which features heavily in the final estimate of Eq. (6), is a completely auxilary quantity. Hypothetically, two different bias distributions could produce the same samples $\left\{\mathbf{x}_{i}\right\}_{i=1}^{B}$, thus producing the same $\left\{f(\mathbf{x}_{i})\right\}_{i=1}^{B}$. But in the calculation of Eq. (6), these two bias distributions would contribute different weights, resulting in different final answers even though the training data from the computer simulator was exactly the same. O’Hagan, (1987) first pointed out this conundrum as a violation of the “Likelihood principle.” A second weakness is the restriction of $q(\mathbf{x})$ to distributional families like GMMs. Failure regions may be complex and may not fit nicely under these target distributions. Consequently, SIIS typically requires large $B$ in order to adequately cover the failure region. For context, Peherstorfer et al., (2016) entertained $B\in\left\{100,1000,10000,100000\right\}$ and found performance continued to improve as $B$ increased. Furthermore, there is no clear avenue for determining the budget allocation between surrogate training and importance sampling estimation. [Throughout, we will borrow from our proposed method to inform budget allocation for the SIIS competitor, which is how we chose to separate the budget of Figure 3 into 100 for training and 50 for estimation.] Finally, the SIIS method does not leverage the full potential of the surrogate model. Many expensive simulator evaluations go into training the surrogate and, in the end, the surrogate is only used to identify predicted failures, with no mention of its UQ. [Notice, as portrayed in the center panel of Figure 3, the black circles are abandoned after the bias distribution is trained.]
3 Two-stage design for failure probability estimation
All in all, we are unsatisfied with the two state-of-the-art surrogate-based methods for estimating $\alpha$ which we just described: surrogate MC and SIIS. While the drawbacks to SIIS are apparent, the limitations of surrogate MC deserve more attention. To investigate further, we conducted the same CL sequential design of the Herbie function (Figure 3), this time storing the surrogate MC estimate $\hat{\alpha}_{\textrm{SURR}}$ (4), which we now denote as simply $\hat{\alpha}$, after every 10th acquisition. Progress is shown in the left panel of Figure 4. While there is significant learning occurring with early acquisitions, it quickly plateaus. By the time the surrogate has been trained on 100 points, it knows just as much about $\alpha$ as it does when it has been trained on the full budget of $N=150$ points. In this case, there is bound to be a more effective way to use the remaining simulation budget. We propose a solution: first developing a stopping criterion to decide when to halt contour location, then proposing a hybrid Monte Carlo estimator that greedily leverages the remaining simulation budget precisely for failure probability estimation.
Figure 4: Left: Surrogate MC failure probability estimate for the Herbie function, updated after every 10th contour locating acquisition. True $\alpha$ in dashed red. Right: Difference in successive estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green. Blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$.
3.1 A stopping criterion for contour location
Let $n$ denote the budget used on contour location and $B$ denote the remaining budget such that the total budget is $N=n+B$. We seek an objective, automatable decision rule for selecting $n$. The right panel of Figure 4 shows the relative difference between each successive $\hat{\alpha}$ estimate from the left panel. As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but they never settle at exactly zero. How small of an update then is small enough to indicate learning has plateaued?
The significance of an update to $\hat{\alpha}$ is only interpretable within the context of $\alpha$, the true failure probability, and $M$, the size of the MC sample used. Small differences might be inconsequential for large $\alpha$ with small $M$, but could be significant for small $\alpha$ with large $M$. To this end, we propose the use of the standard error $\sigma_{\alpha}$ to quantify the relative magnitude of each update. Notice in the estimator of Eq. (2), the indicator $\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$. Thus, the MC estimator has variance inversely proportional to $M$:
$\mathbb{V}[\hat{\alpha}_{\textrm{MC}}]=\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{V}\left[\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}\right]=\frac{\alpha(1-\alpha)}{M}.$
Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator. This standard error intrinsically accounts for the magnitude of $\alpha$ and $M$. Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate.
The green dashed lines in the right panel of Figure 4 show ±$\pm\hat{\sigma}_{\alpha}$, updated along with $\hat{\alpha}$. At the start, updates to $\hat{\alpha}$ are far beyond one standard error. But after a while, the updates sit comfortably within these bounds. Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$. We posit that, for fixed $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is “small enough.” Accordingly, we propose halting CL once we have observed two successive updates within ±$\pm\hat{\sigma}$. The blue square in Figure 4 marks this point, stopping CL for this example at $n=100$, effectively identifying a point where surrogate learning has saturated and leaving a remaining budget of $B=50$ (this informed the budget allocation of Figure 3).
Note, this stopping criterion requires the same $X_{M}$ be used in every $\hat{\alpha}$ calculation. If a new $X_{M}$ is generated for each estimate, there will be additional variability among the $\hat{\alpha}$’s, and the proposed ±$\pm\hat{\sigma}_{\alpha}$ threshold would need to be increased. Removing this extraneous source of variability allows successive estimates to pinpoint differences stemming solely from the surrogate.
Our budgeting scheme is summarized as follows. Initialize a surrogate, begin a contour locating sequential design, and collect one sample $X_{M}$. At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as seen in Figure 4). To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before the contour location could be halted. If the contour location design is succeeding, it will be placing some points in the failure region. We recommend a default of 10 failures, but suggest increasing this for more complex, higher dimensional problems. Once the design has surpassed these minimum requirements, estimate $\hat{\alpha}$ after every 10th acquisition. Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$. When this condition has been satisfied twice in a row, halt the sequential design and proceed to the next stage (the topic of Section 3.2).
Choosing M.
In practice, it is important to select a MC sample size that is large enough without being too cumbersome. Larger $M$ will produce more effective sum approximations of the true integrals, but we need surrogate predictions (e.g., Eq. 3) for every sample in $X_{M}$. Although this computation is significantly cheaper than evaluation of the expensive simulator, it may still be significant for extremely large $M$, particularly with a heftier surrogate like a Bayesian DGP. If we knew $\alpha$, we could determine $M$ by thresholding the standard error $\sigma_{\alpha}$. For example, to achieve a standard error less than 10% of $\alpha$ requires:
To achieve this precision, $\alpha=10^{-4}$ requires $M$ nearly 1 million, $\alpha=10^{-5}$ requires $M$ nearly 10 million, and $\alpha=10^{-6}$ requires $M$ nearly 100 million. In practice, with unknown $\alpha$, we recommend either choosing $M$ conservatively given expert insight into the anticipated failure probability or performing a pilot study where $M$ is incremented say by one million at a time. Gather surrogate predictions along this incremental process, and use $\hat{\alpha}$ to inform $\hat{\sigma}_{\alpha}$.
3.2 Hybrid Monte Carlo for failure probability estimation
Once $n$ is determined, how should we spend the remaining budget of simulator evaluations ($B=N-n$), in conjunction with our trained surrogate, to most effectively estimate $\alpha$? We propose an upgrade to the surrogate MC estimator of Eq. (4). Let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive). Then, Eq. (1) may be partitioned as
@@ 3.2 Hybrid Monte Carlo for failure probability estimation
inspiring an equivalent formulation of the surrogate MC estimator,
@@
Now for $\mathbf{x}_{i}\in\mathcal{U}$, upgrade surrogate prediction $\mu_{n}(\mathbf{x}_{i})$ to expensive evaluation of the true $f(\mathbf{x}_{i})$, resulting in the following “hybrid MC” estimator:
[…]
The hybrid nature of this estimator is in the use of both true simulator evaluations (left term) and surrogate evaluations (right term). The accuracy of this estimator is contingent upon the classification accuracy of the surrogate and the choice of $\mathcal{U}$. As long as we restrict the size of $\mathcal{U}$ so only $B$-many samples from $X_{M}$ fall within it, i.e., $\sum_{i=1}^{M}\mathbbm{1}_{\left\{\mathbf{x}_{i}\in\mathcal{U}\right\}}=B$, then we will meet our budgetary constraints.
@@
Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize in the context of the discrete set $X_{M}$. In a perfect world, $\mathcal{U}$ would capture all the samples which the surrogate would misclassify. We are not privy to this information, and we are under a limited budget, but we can leverage the surrogate’s UQ to identify the $B$-many samples at which it is most uncertain. Recall, classification entropy (5) quantifies the degree of uncertainty in the pass/fail prediction of a surrogate. Consequently, our proposed allocation of samples to $\mathcal{U}$ is strikingly simple – place the $B$-many samples from $X_{M}$ with the highest entropy into $\mathcal{U}$. Denote these as $X_{M}^{B}$. Revisiting Figure 2, notice the high entropy regions (green dotted/dashed) align well with the misclassification regions (yellow shading). The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations to correct inaccurate surrogate predictions.
Ultimately, after evaluating $f(\mathbf{x})$ for $\mathbf{x}\in X_{M}^{B}$, we update the surrogate with these additional observations, thereby updating $\mu_{n}(\mathbf{x})$, before returning $\hat{\alpha}_{\textrm{HYB}}$ (7). Since the $X_{M}^{B}$ points are likely near a failure contour, they could significantly improve the accuracy of the surrogate predictions for other samples near the contour. For reference, consider the right panel of Figure 3. The black circles showing the contour locating design for the GP surrogate are unchanged, but here the additional budget of 50 evaluations is spent on the MC samples from $p(\mathbf{x})$ with the highest classification entropy (blue triangles). Compared to the left panel which continues to explore the entire domain, these highest entropy samples exclusively exploit the predicted failure contour. Compared to the bias samples portrayed in the middle panel, where the original surrogate design is abandoned, our two-stage design is able to leverage all simulator evaluations in its final estimate.
3.3 Summary and related work
We have presented a comprehensive strategy for estimating Eq. (1) from a budget of size $N$. It is summarized as follows. Start with a very large Monte Carlo sample from $p(\mathbf{x})$, denoted $X_{M}$. Train a surrogate model through a contour locating sequential design, monitoring the progress of the $\hat{\alpha}_{\textrm{SURR}}$ estimate for given $X_{M}$, and halting contour location once two consecutive updates are within ±$\pm\hat{\sigma}_{\alpha}$. Denote the design points chosen in this stage as $X_{n}$. Then, use the trained surrogate to identify the highest entropy samples, $X_{M}^{B}$, evaluate the simulator at these inputs, and update the surrogate with these new observations. Finally, return $\hat{\alpha}_{\textrm{HYB}}$ following Eq. (7).
Given the deterministic nature of $f(\mathbf{x})$, effective surrogates should interpolate observed points. In this light, the hybrid MC estimator (7) is akin to a surrogate MC estimator (4) when the surrogate has been trained on all $\mathbf{x}_{i}\in\mathcal{U}$. This connection permits us to frame our contribution as a two-stage design. In the first stage, $X_{n}$ is chosen over $\mathcal{X}$ based on contour location. In the second stage, $X_{M}^{B}$ is chosen over $\mathbbm{P}_{\mathbf{x}}$ based on maximum entropy. The final design of size $N$ is then $X_{N}=[X_{n}^{\top},{X_{M}^{B}}^{\top}]^{\top}$.